Find the refractive index of the material of a prism of angle 59°42', if the angle of minimum deviation produced for a perticular colour of light is 39°28'.
Also find angle of incidence
1
Expert's answer
2020-09-06T17:22:08-0400

First convert radians to degrees. Thus:


"59\degree 42= 59+ \frac {42}{60}= 59.7\degree"


"39\degree 42= 39+ \frac {28} {60}=39.46\degree"


"n=\sin \frac {59.7+39.46}{2}= 0.761"


"\sin \frac {59.7} {2}= 0.497"


Thus refractive index we divide the two: "\frac {0.761}{0.497}= 1.53"

"n=1.53"


"\delta_m= 2i-A"


Therefore: "39.46\degree= 2i-59.7\degree"


"i=49.58\degree"